# Equal contribution    Project lead    Corresponding author

### v1

@@ Appendix C Processed Data Archive
In some data collection systems, missing values are often filled with 00 or another constant, leading to sequences with constant values that do not represent valid patterns for the model. To address this, we developed a filtering method that uses a fixed-length window to scan the entire sequence. For each window, we calculate the ratio of first-order to second-order differences, discarding the window if this ratio exceeds a pre-specified threshold (set to 0.2 in our case). The remaining valid window sequences are then concatenated into a single sequence. This process transforms the original sequence into multiple sub-sequences, effectively removing segments with invalid patterns.
Following the processing steps described above, we compiled a high-quality time series dataset named Time-300B, which spans a range of sampling frequencies from seconds to yearly intervals, encompassing a total of 309.09 billion time points. To optimize memory efficiency and loading speed, each dataset is split into multiple binary files, with a metafile providing details such as the start and end positions of each sequence. This setup allows us to load the data using a fixed amount of memory during training, preventing memory shortages. Datasets like Weatherbench, CMIP6, and ERA5 are particularly large, often leading to data imbalance and homogenization. To mitigate these issues, we apply down-sampling to these datasets. During training, we utilized approximately 117 billion time points, sampling each batch according to fixed proportions of domains and distributions of observation values.
Below, we outline the key properties of the datasets after processing, including their domain, sampling frequency, number of time series, total number of observations, and data source.
@@
Weatherbench (Hourly) Nature H 3,984,029 74,630,250,518 Ansari et al. (2024)
Weatherbench (Daily) Nature D 301,229 3,223,513,345 Ansari et al. (2024)
Weatherbench (Weekly) Nature W 226,533 462,956,049 Ansari et al. (2024)
@@ D.1 Ablation Study
ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Average
Time-MoEbase 0.357 0.305 0.338 0.201 0.160 0.113 0.245
         w/o Huber loss 0.365 0.309 0.344 0.205 0.163 0.115 0.250
         w/o multi-resolution layer 0.358 0.313 0.348 0.212 0.164 0.116 0.252
         w/o mixture-of-experts 0.370 0.317 0.347 0.212 0.163 0.121 0.255
         w/o auxiliary loss 0.368 0.325 0.350 0.219 0.164 0.123 0.258
As shown in Table 11, replacing the MoE layers with standard FFNs (denoted as “w/o mixture-of-experts ”) led to a noticeable performance decline, with the MSE worsening from $0.245$ to $0.255$. This highlights the significant contribution of the sparse architecture to the model’s overall performance, as its dynamic routing enables more specialized processing of diverse input patterns.
We also conducted experiments by retaining only the horizon-32 forecasting head from the Time-MoEbase (denoted as “w/o multi-resolution layer”), excluding the multi-task optimization. The performance of this modified model was slightly inferior to the complete Time-MoEbase, highlighting the importance of multi-resolution outputs in capturing various temporal dependencies.
@@
ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Average MSE Inference Speed
Time-MoEbase 0.357 0.305 0.338 0.201 0.160 0.113 0.245 0.095 s/iter
Time-MoEbase w/ {1,8,32} 0.353 0.316 0.370 0.225 0.161 0.112 0.256 0.130 s/iter
Time-MoEbase w/ {1,8} 0.389 0.391 0.441 0.304 0.174 0.117 0.302 0.411 s/iter
Time-MoEbase w/ {1} 1.071 0.920 2.098 2.320 1.500 0.171 1.346 2.834 s/iter
@@ D.2 Training Precision Analysis
ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Average MSE Training Speed Inference Speed Training Memory Inference Memory
Time-MoEbase 0.357 0.305 0.338 0.201 0.160 0.113 0.245 0.84 s/iter 0.095 s/iter 1.77 GB 226.70 MB
Time-MoEbase w/o FA 0.357 0.305 0.338 0.201 0.160 0.113 0.245 1.09 s/iter 0.118 s/iter 1.77 GB 226.70 MB
Time-MoEbase w/ FP32 0.358 0.303 0.342 0.198 0.158 0.110 0.244 1.24 s/iter 0.133 s/iter 2.21 GB 453.41 MB
@@ Appendix E Forecast Showcases
To visualize the performance differences among various large-scale time series models, we present the forecasting results of our model, Time-MoE, in comparison to the ground truth across six real-world benchmarks. These benchmarks include ETTh1, ETTh2, ETTm1, ETTm2, Weather, and Electricity datasets. Alongside Time-MoE’s results, we also show the performance of other large-scale baseline models at different scales, providing a comprehensive view of their comparative capabilities (Figures 510). In all figures, the context length is set to 512, and the forecast horizon is 96. To enhance clarity and aesthetics, we display the full forecast output, complemented by a portion of the preceding historical input data, ensuring a more intuitive comparison.
@@
Figure 10: Zero-shot forecasting cases from Electricity by different models, with forecast horizon 96.

### v2

@@ Appendix C Processed Data Archive
In some data collection systems, missing values are often filled with 00 or another constant, leading to sequences with constant values that do not represent valid patterns for the model. To address this, we developed a filtering method that uses a fixed-length window to scan the entire sequence. For each window, we calculate the ratio of first-order and second-order differences, discarding the window if this ratio exceeds a pre-specified threshold (set to 0.2 in our case). The remaining valid continuous window sequences are then concatenated into a single sequence. This process transforms the original sequence into multiple sub-sequences, effectively removing segments with invalid patterns.
Following the processing steps described above, we compiled a high-quality time series dataset named Time-300B, which spans a range of sampling frequencies from seconds to yearly intervals, encompassing a total of 309.09 billion time points. To optimize memory efficiency and loading speed, each dataset is split into multiple binary files, with a metafile providing details such as the start and end positions of each sequence. This setup allows us to load the data using a fixed amount of memory during training, preventing memory shortages. Datasets like Weatherbench, CMIP6, and ERA5 are particularly large, often leading to data imbalance and homogenization. To mitigate these issues, we apply down-sampling to these datasets. During training, we utilized approximately 117 billion time points in Time-300B, sampling each batch according to fixed proportions of domains and distributions of observation values.
Below, we outline the key properties of the datasets after processing, including their domain, sampling frequency, number of time series, total number of observations, and data source. Also, we present the KEY COMPONENT’S SOURCE CODE of the data-cleaning pipeline in Algorithm 2.
@@
Weatherbench (Hourly) Nature H 3,984,029 74,630,250,518 Rasp et al. (2020)
Weatherbench (Daily) Nature D 301,229 3,223,513,345 Rasp et al. (2020)
Weatherbench (Weekly) Nature W 226,533 462,956,049 Rasp et al. (2020)
@@
Algorithm 2 Sample code of Data-cleaning Pipline
# Missing Value Processing
def split_seq_by_nan_inf(seq, minimum_seq_length: int = 1):
output = []
sublist = []
for num in seq:
if num is None or np.isnan(num) or np.isinf(num):
if len(sublist) >= minimum_seq_length:
output.append(sublist)
sublist = []
else:
sublist.append(num)
if len(sublist) >= minimum_seq_length:
output.append(sublist)
return output
# Invalid Observation Processing
def split_seq_by_window_quality(seq, window_size: int = 128, zero_threshold, minimum_seq_length: int = 256):
if len(seq) <= window_size:
flag, info = check_sequence(seq, zero_threshold=zero_threshold)
if flag:
return [seq]
else:
return []
i = window_size
sub_seq = []
out_list = []
while True:
if i + window_size > len(seq):
window_seq = seq[i - window_size: len(seq)]
i = len(seq)
else:
window_seq = seq[i - window_size: i]
flag, info = check_sequence(window_seq, zero_threshold=zero_threshold)
if flag:
sub_seq.extend(window_seq)
else:
if len(sub_seq) >= minimum_seq_length:
out_list.append(sub_seq)
sub_seq = []
if i >= len(seq):
break
i += window_size
if len(sub_seq) >= minimum_seq_length:
out_list.append(sub_seq)
return out_list
def check_sequence(seq, zero_threshold: float):
import numpy as np
if not isinstance(seq, np.ndarray):
seq = np.array(seq)
if len(seq.shape) > 1:
raise RuntimeError(f’Dimensionoftheseqisnotequalto1:{seq.shape}’)
flag = True
info = {}
nan_count = np.sum(np.isnan(seq))
info[’nan_count’] = nan_count
if nan_count > 0:
flag = False
return flag, info
inf_count = np.sum(np.isinf(seq))
info[’inf_count’] = inf_count
if inf_count > 0:
flag = False
return flag, info
zero_ratio = np.sum(seq == 0) / len(seq)
info[’zero_ratio’] = zero_ratio
if zero_ratio > zero_threshold:
flag = False
first_diff = seq[1:] - seq[:-1]
first_diff_zero_ratio = np.sum(first_diff == 0) / len(first_diff)
info[’first_diff_zero_ratio’] = first_diff_zero_ratio
if first_diff_zero_ratio > zero_threshold:
flag = False
second_diff = seq[2:] - seq[:-2]
second_diff_zero_ratio = np.sum(second_diff == 0) / len(second_diff)
info[’second_diff_zero_ratio’] = second_diff_zero_ratio
if second_diff_zero_ratio > zero_threshold:
flag = False
return flag, info
@@ D.1 Ablation Study
ETTh1 ETTh2 ETTm1 ETTm2 Weather Global Temp Average
Time-MoEbase 0.357 0.305 0.338 0.201 0.160 0.211 0.262
         w/o Huber loss 0.365 0.309 0.344 0.205 0.163 0.217 0.267
         w/o multi-resolution layer 0.358 0.313 0.348 0.212 0.164 0.217 0.269
         w/o mixture-of-experts 0.370 0.317 0.347 0.212 0.163 0.223 0.272
         w/o auxiliary loss 0.368 0.325 0.350 0.219 0.164 0.226 0.275
As shown in Table 11, replacing the MoE layers with standard FFNs (denoted as “w/o mixture-of-experts ”) led to a noticeable performance decline, with the MSE worsening from $0.262$ to $0.272$. This highlights the significant contribution of the sparse architecture to the model’s overall performance, as its dynamic routing enables more specialized processing of diverse input patterns.
We also conducted experiments by retaining only the horizon-32 forecasting head from the Time-MoEbase (denoted as “w/o multi-resolution layer”), excluding the multi-task optimization. The performance of this modified model was slightly inferior to the complete Time-MoEbase.
@@
ETTh1 ETTh2 ETTm1 ETTm2 Weather Global Temp Average MSE Inference Speed
Time-MoEbase 0.357 0.305 0.338 0.201 0.160 0.211 0.262 0.095 s/iter
Time-MoEbase w/ {1,8,32} 0.353 0.316 0.370 0.225 0.161 0.213 0.273 0.130 s/iter
Time-MoEbase w/ {1,8} 0.389 0.391 0.441 0.304 0.174 0.222 0.320 0.411 s/iter
Time-MoEbase w/ {1} 1.071 0.920 2.098 2.320 1.500 0.383 1.382 2.834 s/iter
@@ D.2 Training Precision Analysis
ETTh1 ETTh2 ETTm1 ETTm2 Weather Global Temp Average MSE Training Speed Inference Speed Training Memory Inference Memory
Time-MoEbase 0.357 0.305 0.338 0.201 0.160 0.211 0.262 0.84 s/iter 0.095 s/iter 1.77 GB 226.70 MB
Time-MoEbase w/o FA 0.357 0.305 0.338 0.201 0.160 0.211 0.262 1.09 s/iter 0.118 s/iter 1.77 GB 226.70 MB
Time-MoEbase w/ FP32 0.358 0.303 0.342 0.198 0.158 0.208 0.261 1.24 s/iter 0.133 s/iter 2.21 GB 453.41 MB
@@ Appendix E Forecast Showcases
To visualize the performance differences among various large-scale time series models, we present the forecasting results of our model, Time-MoE, in comparison to the ground truth across six real-world benchmarks. These benchmarks include ETTh1, ETTh2, ETTm1, ETTm2, Weather, and Global Temp datasets. Alongside Time-MoE’s results, we also show the performance of other large-scale baseline models at different scales, providing a comprehensive view of their comparative capabilities (Figures 510). In all figures, the context length is set to 512, and the forecast horizon is 96. To enhance clarity and aesthetics, we display the full forecast output, complemented by a portion of the preceding historical input data, ensuring a more intuitive comparison.
@@
Figure 10: Zero-shot forecasting cases from Global Temp by different models, with forecast horizon 96.